OpenStudy (anonymous):

@iGreen @cwrw238 help meeee plzzz

OpenStudy (anonymous):

whats the question?

OpenStudy (anonymous):

You are decorating a cylindrical hat box to give to your grandmother as a gift. You want to paste family photos onto the curved part of the box. The height of the box is 8 inches and the radius is 6 inches. If none of the photos overlap, what is the area of photographs you need? Use three point one four for pi . one hundred fifty point seven two square inches three hundred one point four four square inches nine hundred four point three two square inches one thousand two hundred five point seven six square inches

OpenStudy (anonymous):

i got 527

OpenStudy (anonymous):

it is not one of the answer choices

OpenStudy (igreen):

You're looking for the Surface Area of a Cylinder. Formula: \(2 \pi r^2 + 2 \pi r h\) Plug in what we know: \((2)(3.14)(6^2) + (2)(3.14)(6)(8)\) Simplify Exponent: \((2)(3.14)(36) + (2)(3.14)(6)(8)\) Okay, can you finish simplifying that? @_lilb_to_smooth_4_U_

OpenStudy (anonymous):

ill try wait don't leave yet @iGreen

OpenStudy (anonymous):

multiply

OpenStudy (igreen):

Yes..

OpenStudy (igreen):

Hmm..something's wrong..I don't know. I come up with 527 also..

OpenStudy (igreen):

Hold on.

OpenStudy (anonymous):

i know is it a typo

OpenStudy (igreen):

What do you mean?

OpenStudy (igreen):

Ohhh, I get it now..it doesn't want the area of the tops, just the area of the side. \(2 \pi r h\) Plug in what we know: \((2)(3.14)(6)(8)\) Can you multiply all of those?

OpenStudy (igreen):

@_lilb_to_smooth_4_U_

OpenStudy (anonymous):

301.44

OpenStudy (igreen):

Yep, that's your answer.